OpenStudy (anonymous):

calculus help...plz explain and answer question...i totally dont know what they are...studying on my own

OpenStudy (phi):

r is a position vector... the position of a point that changes with time. it has components in the x direction (the "ith" component) and y direction the velocity of the point is dr/dt (derivative with respect to time) you take the derivative of each component separately.... you get a new vector that represents the velocity in the x and y dimensions...

OpenStudy (phi):

it is not obvious, but they mean v is the velocity vector and a is the acceleration vector (2nd derivative of r, and the 1st derivative of v)

OpenStudy (anonymous):

(t-sint)i = (1-cost)i = (sint)i (1-cost)j = (sint)j = (cost)j f(x) f'(x) f''(x)

OpenStudy (anonymous):

now??

OpenStudy (phi):

r= (t-sin t) i + (1-cos t) j v= dr/dt = (1-cos t) i + sin t j or v= < (1-cos t, sin t> the magnitude of v is found by doing the dot product |v|^2 = v dot v and |v| = sqrt( v dot v)

OpenStudy (anonymous):

what is dot product

OpenStudy (phi):

Khan has videos on vectors. See http://www.khanacademy.org/math/linear-algebra/vectors_and_spaces/dot_cross_products/v/vector-dot-product-and-vector-length Here is one from the physics playlist http://www.khanacademy.org/science/physics/electricity-and-magnetism/v/the-dot-product which might be better. But the dot product is the sum of the product of corresponding elements. In this case < (1-cos t, sin t> dot < (1-cos t, sin t> = (1-cos t)^2 + (sin t)^2 which simplifies to |v| = 2(1- cos t)

OpenStudy (anonymous):

ohk got it...so how to get minimum nd maximum values

OpenStudy (phi):

we know cos t has a max of +1 and a min of -1, so |v| ranges between 2(1-1) and 2(1 - (-2)) or 0 and 4

OpenStudy (anonymous):

agree

OpenStudy (phi):

** 2(1 - (-1)) or 0 and 4

OpenStudy (anonymous):

plz carry on

OpenStudy (phi):

a= < sin t , cos t> can you find its dot product (that gives you |a|^2 )

OpenStudy (anonymous):

yes it is 1

OpenStudy (anonymous):

sin^2t+cos^2t = 1

OpenStudy (anonymous):

1^2 = 1

OpenStudy (anonymous):

am i on track?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok so min and max for velocity is 0 and 4 repectively and min and max for acceleration is 1 and 1...correct?

OpenStudy (phi):

the magnitude of the acceleration is always 1, but its direction is changing.

OpenStudy (anonymous):

oh ya but we do not need to care about direction...right?

OpenStudy (phi):

you might care if you were trying to analyze what is going on... but you don't need the direction if all you care about is the magnitude.

OpenStudy (anonymous):

ok so how to do rest of the part

OpenStudy (phi):

I am not sure what they want for a plot... for the smaller hoop, make the frequency twice as fast. that means in sin t and cos t replace t with 2t this will make the hoop spin twice as fast.

OpenStudy (anonymous):

ok so what is the vector function that models the behavior